You are considering taking a short position in one-year futures on a stock. The stock is expected to pay two dividends; each dividend is $4 per share; first dividend is to be paid in four months, and the second dividend is paid in ten months. The current stock price is $100, and the risk-free rate of interest is 10% for all maturities. What should be one-year futures price contract?
Answer #1

Future Price, F0 = (S0-I)erT

Here, S0 = Spot price = $100

r= risk free rate = 10%

T= time period = 1 year

Calculating I = Present value of dividends paid

= 4/ (1+(10%/12))4 + 4/(1+(10%/12))10

=3.87+3.68

=7.55

Thus, F0 = (100-7.55)*e(0.10*1)

Future Price, F0 = 102.17